ABC Company has set the following standards in order to produce one unit
of its single product:

                    standard quantity           standard price
direct materials     2.7 yards                  ?? per yard
direct labor         2 hours                    $14 per hour
variable overhead    2 hours                    ?? per hour

During August, ABC Company spent $133,770 to purchase direct materials and
had direct labor totaling $189,440. During August, ABC Company used 16,000
yards of direct materials in the production of units. ABC had 2,500 yards
of materials in its August 1 direct material inventory and had 1,200 yards
yards remaining in its direct materials inventory at August 31.

ABC Company reported the following variances for August:

  Direct material price variance ..............  $12,495 unfavorable
  Direct labor rate variance ..................  $24,240 unfavorable
  Total direct labor variance .................  $6,560 favorable
  Variable overhead spending variance .........  $1,900 unfavorable
  Variable overhead efficiency variance .......  $20,900 favorable

Calculate ABC's direct material quantity variance for August. If the
variance is favorable, place a minus sign in front of your answer
(i.e., -5000). If the variance is unfavorable, simply enter your
answer as a number (i.e., 5000).
Answer #1

Direct material purchases in yards = 16,000+1,200-2,500 = 14,700 yards

Actual price of direct material per yard = $133,770/14,700 = $9.1 per yard

Direct material price variance = Actual quantity * (Standard price-Actual price)

Let standard price per yard be x

-$12,495 = 14,700 * ($x - $9.1)

-12,495 = $14,700 x - 133,770

-$14,700 x = -$133,770+12,495

$14,700 x = $121,275

x = $121,275 / 14,700 = $8.25

Hence Standard price per yard is $8.25

Direct material quantity variance = Standard price * (Standard quantity-Actual quantity)

Standard quantity = 7,000*2.7 = 18,900 yards

Direct material quantity variance = $8.25 * (18,900-16,000) = -23,925 Favorable

Working notes:

Total direct labor variance = Standard hours*Standard rate - Actual hours*Actual rate

$6,560 F = Standard hours*Standard rate - $189,440

Standard hours*Standard rate = $189,440 + 6,560 = $196,000

Let standard hours be x

x * $14 = $196,000

x = $196,000 / 14 = 14,000 hours

Hence standard hours = 14,000

Actual units produced = 14,000 / 2 = 7,000 units